OpenStudy (anonymous):

∑(3n/√(n^3-1) use ratio test to find weather series convergent or divergent ?

OpenStudy (anonymous):

the ratio test is when you take : an + 1 / an for you condition : \[ an= 3n/ \sqrt{n^3-1}\] so: \[an + 1/ an = (3n + 1/ \sqrt{n^3-1 +1} )/( 3n/\sqrt{n^3-1})\] simplify ^_^

OpenStudy (anonymous):

then take the limit as x approaches infinity :)

OpenStudy (anonymous):

(n-1)^3 is not equal to n^3-1^3 just review your solution